OpenStudy (anonymous):

log(w-3)+log(w)=log(28) log is in base 10

OpenStudy (anonymous):

solve for w

OpenStudy (ghazi):

log a +log b= log (ab)

OpenStudy (ghazi):

log (w-3)+log w= log ((w-3)w)=log28 therefore w(w-3)=28...can you solve this?

OpenStudy (anonymous):

quadratic formula?

OpenStudy (anonymous):

w^2-3w-28=0

OpenStudy (ghazi):

yes solve it

OpenStudy (anonymous):

so its 7 and not -4 right?

OpenStudy (ghazi):

w^2-7w+4w-28=0

OpenStudy (anonymous):

what? where do you get 7w and 4w?

OpenStudy (ghazi):

yes it's 7...coz for negative number log is not defined

OpenStudy (ghazi):

@Ldaniel clear?

OpenStudy (anonymous):

no

OpenStudy (ghazi):

why ?

OpenStudy (anonymous):

oh wait cause 7-4 =3

OpenStudy (ghazi):

yes i split -7w+4w=-3w